OpenStudy (anonymous):

A wagon train that is one mile long advances one mile at a constant rate. During the same time period, the wagon master rides his horse at a constant rate from the front of the wagon train to the rear, and then back to the front. How far did that wagon master ride? My teacher hates me its 10 extra credit points...anyone?

OpenStudy (anonymous):

a constant rate but you do not know the constant rate?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

well the trick answer is two miles, since the train is one mile long and the "wagon master" rode from the front to the back and then to the front but i am not sure if this is the answer you need

OpenStudy (anonymous):

oh i see nvm that is wrong

OpenStudy (anonymous):

lol i don't get it either

OpenStudy (anonymous):

we can do this, give me a minute

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

ok give me another few minutes

OpenStudy (anonymous):

lol its takin me forever

OpenStudy (anonymous):

think i am almost there

OpenStudy (anonymous):

i think its impossible

OpenStudy (anonymous):

no no less than a minute now

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

let me know when you are ready, i will write a solution

OpenStudy (anonymous):

first of all since it makes no difference apparently how fast the train goes, lets make it easy and make it go one mile per hour, so it completes its journey in one hour

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and what we are going to solve for is the rate of the horse, because if we know the rate of the horse we know the total distance. they will be the same since the total time will be one hour

OpenStudy (anonymous):

i put the rate of the horse as \(x\) and the rate of the train is 1, so as the horse travels to the rear of the train, the combined rate is \(1+x\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok the horse travels one mile to the rear of the train at a speed of \(1+x\) the time taken to reach the end of the train is distance divided by rate, so it takes \[\frac{1}{1+x}\] to reach the back

OpenStudy (anonymous):

now the horse turns around and heads to the front of the train. since it is going with the train the combined rate is now \(x-1\) and it has to complete its one mile journey at then end of one hour it has already used up \(\frac{1}{1+x}\) hours leaving \[1-\frac{1}{x+1}\] hours to go the one mile

OpenStudy (anonymous):

using distance equals rate times time, we get \[1=(x-1)(1-\frac{1}{x+1})\]

OpenStudy (anonymous):

let me know if you are still with me, because this is really the end. just a matter of solving this equation

OpenStudy (radar):

@satellite73 , I got x=3, I think that is the speed of the horse ?? 3 mph

OpenStudy (anonymous):

i got \(1+\sqrt{2}\) this took me quite a while then i googled and got an answer i didn't understand, but it was still \(1+\sqrt{2}\)

OpenStudy (anonymous):

i mean i didn't understand the method used to solve the problem. i am sure it was fine, but i did something else

OpenStudy (radar):

O.K I am sure the speed of the wagon master was the same as the horse, that is about all that I am sure of. lol

OpenStudy (anonymous):

truth be told i wolframed to solve the equation. i guess i should do it with pencil and paper

OpenStudy (anonymous):

lol i got 3 too, but we both must have made the same mistake!

OpenStudy (anonymous):

algebra challenged today

OpenStudy (radar):

Well I worked it differently and got \[1\pm \sqrt{2}\] I think I will go get a cup of coffee, Have a nice day satellite73.

OpenStudy (radar):

Just checked your link, I got close lol

OpenStudy (anonymous):

second time i got \[1=x-1-\frac{x-1}{x+1}\] \[\frac{x-1}{x+1}=x-2\] \[x^2-2x-1=0\] so now i am happy yes, it is coffee time have a good one